OpenStudy (taylor.joe):

Help please! When will the equation xn = c, where c is a whole number, have two solutions?

OpenStudy (daniel.ohearn1):

When do equations in general have two solutions for one variable?

OpenStudy (daniel.ohearn1):

Do you about the quadratic function yet?

OpenStudy (taylor.joe):

Yes. I do.

OpenStudy (taylor.joe):

Because it can be positive or negative?

OpenStudy (daniel.ohearn1):

Yes and why is the solution positive or negative?

OpenStudy (taylor.joe):

That i'm unsure about. Its been a long time since i've actually done these kind of problems.

OpenStudy (daniel.ohearn1):

It's because of the discriminant, and the fact that x^2 = (-x)^2

OpenStudy (taylor.joe):

Ohhhh. Okay. That makes sense.

OpenStudy (daniel.ohearn1):

The discriminant is the part of the quadratic that discrimates between unique solutions so what part do you know that to be?

OpenStudy (daniel.ohearn1):

x^2 = (-x)^2 is the clue

OpenStudy (taylor.joe):

I'm sorry. I really don't understand this stuff. Also i'm a slow learner.

OpenStudy (daniel.ohearn1):

http://www.purplemath.com/modules/solvquad4.htm

OpenStudy (daniel.ohearn1):

This might help

OpenStudy (taylor.joe):

Thank you! I appreciate it. Sorry for any inconveniences.

OpenStudy (daniel.ohearn1):

The thing to understand about this is how we can two solutions, one positive and one negative from an equation..

OpenStudy (taylor.joe):

Okay. Thank you. :)

OpenStudy (daniel.ohearn1):

We can't just multiply by a number, right, how could a function say y=4x have two solutions?

OpenStudy (daniel.ohearn1):

two solutions for x?

OpenStudy (daniel.ohearn1):

it would appear there is only one x = y/4 right

OpenStudy (daniel.ohearn1):

?

OpenStudy (taylor.joe):

Yes?

OpenStudy (daniel.ohearn1):

Turns out if we want two solutions, we can essentially change the magnitude of the function.. So we have x^2 = (y/4)^2 , then how would u solve for x?

OpenStudy (daniel.ohearn1):

What do we do to undo squaring a function?

OpenStudy (taylor.joe):

Square root it?

OpenStudy (daniel.ohearn1):

Yes so you would get the sqrt of x^2 = the sqrt of y^2 over the sqrt of 4^2

OpenStudy (daniel.ohearn1):

What is the sqrt of x^2?

OpenStudy (taylor.joe):

Would it be just x?

OpenStudy (daniel.ohearn1):

Right then what is the sqrt of 16 though?

OpenStudy (taylor.joe):

4 right?

OpenStudy (daniel.ohearn1):

remember what I said before now x^2 = (-x)^2

OpenStudy (taylor.joe):

Yeah

OpenStudy (daniel.ohearn1):

So 16= x^2 = (-x)^2 The sqrt of 16 is therefore not only 4

OpenStudy (taylor.joe):

so its 4 and -4?

OpenStudy (daniel.ohearn1):

So we have two solutions then right? If y=4x , x = sqrt y^2 / sqrt 4 therefore x= y / 2 or x= y/-2

OpenStudy (taylor.joe):

I'm confused on what you just said. Im sorry.

OpenStudy (daniel.ohearn1):

When you take the square root of a fraction by the way you'd apply the square root to both numerator and the denominator

OpenStudy (taylor.joe):

Okay. So what's the question again?

OpenStudy (daniel.ohearn1):

In the problem you're dealing with now our y is just c, a whole number, like 1,2,3,4,5

OpenStudy (taylor.joe):

Okayyy. So the question i asked basically is asking when will there be two solutions for c?

OpenStudy (daniel.ohearn1):

c also stands for constant many cases in math.. Say it's 4 4= nx

OpenStudy (daniel.ohearn1):

The solutions are traditionally in algebra for first of all x

OpenStudy (daniel.ohearn1):

So the question is what would n have to be in order for there to be 2 solutions for x?

OpenStudy (daniel.ohearn1):

Also in most math books n is an integer, meaning it's a number we would count on a number line but it can be positive or negative

OpenStudy (daniel.ohearn1):

n could be -5, -4,-3,-2,-1,0, 1, 2, 3, 4, 5, 6, 7, 8... or so on

OpenStudy (daniel.ohearn1):

Putting everything together we've been talking about what do you think n would have to be as a general rule?

OpenStudy (taylor.joe):

I'm still confused.

OpenStudy (daniel.ohearn1):

(-x)^2 = (x)^2 sqrt of 16 is 4 and -4

OpenStudy (taylor.joe):

Isn't it when x is either a positive or a negative?

OpenStudy (daniel.ohearn1):

Yes but why is it positive or negative?

OpenStudy (taylor.joe):

Because it would make the answer either a positive or negative?

OpenStudy (daniel.ohearn1):

How does x become positive or negative if c=nx What is n?

OpenStudy (daniel.ohearn1):

it's not a positive and a negative at the same time it's a positive or it's a negative in what sort of situation?

OpenStudy (daniel.ohearn1):

You said it before

OpenStudy (taylor.joe):

Would it have to do with it being squared or sqrt?

OpenStudy (daniel.ohearn1):

Yes now its asking how do we get two solutions if x= (-2)^2= 4 we have one solution but if we have sqrt of 4

OpenStudy (taylor.joe):

do you use the quadratic formula for any of this?

OpenStudy (daniel.ohearn1):

Well the quadratic tells us -b +or- the sqrt of a discriminant all over 2 will give your solutions

OpenStudy (taylor.joe):

Oh okay.

OpenStudy (daniel.ohearn1):

If the discriminant we were taking the sqrt of was 5 how many solutions would we get?

OpenStudy (daniel.ohearn1):

There would be two right?

OpenStudy (taylor.joe):

i think so one would be negative and one would be positive right?

OpenStudy (daniel.ohearn1):

There will be two solutions thanks to our discriminant

OpenStudy (daniel.ohearn1):

Yes, one discriminant would be negative and one positive

OpenStudy (taylor.joe):

So what does n have to do with all of this?

OpenStudy (daniel.ohearn1):

What does this tell us about n if x=y/n

OpenStudy (taylor.joe):

n is the discriminant ?

OpenStudy (daniel.ohearn1):

n is in the form of a discriminant but it isn't the discriminant because for this problem it is not necessary to use the quadratic formula

OpenStudy (daniel.ohearn1):

what is the discriminant in form of?

OpenStudy (taylor.joe):

I apologize something came up.

OpenStudy (daniel.ohearn1):

That makes it yield two solutions?

OpenStudy (daniel.ohearn1):

I'm trying not to give you the answer but you are so incredibly close to it

Directrix (directrix):

If the problem is this: x^n = c, where c is a whole number, then let n = 2 if you want 2 solutions. x^2 = c x = √c or x = -√c @Taylor.Joe

Directrix (directrix):

Read up on the Fundamental Theorem of Algebra here: https://www.mathsisfun.com/algebra/fundamental-theorem-algebra.html

OpenStudy (taylor.joe):

Gosh. Im so sorry again.

OpenStudy (taylor.joe):

So what's the question again?

OpenStudy (daniel.ohearn1):

In this case, c being a whole number = n x what is n to yield two solutions? You answered it back when you acknowledged there being two solutions for sqrt of 16

OpenStudy (daniel.ohearn1):

what is n in order for you to get two solutions?

OpenStudy (taylor.joe):

Im so sorry dude. its x^n not xn

OpenStudy (taylor.joe):

Im so sorry if you wasted your time. I feel really bad.

OpenStudy (daniel.ohearn1):

so c= x^n ? Well that should be even easier

OpenStudy (taylor.joe):

Okay. Can you please explain that for me?

OpenStudy (daniel.ohearn1):

What do we know about x^2 that's so unique?

OpenStudy (taylor.joe):

I will be right back and answer the question. I have to go some where real quick. I reallly appreiate your help.

OpenStudy (daniel.ohearn1):

Do you have a graphing calculator?

OpenStudy (taylor.joe):

Yes

OpenStudy (daniel.ohearn1):

If you graph that thing, you'll see at point where y= 4 you will have both x=-2 and x=2

OpenStudy (daniel.ohearn1):

cause it's graph looks like a U

OpenStudy (daniel.ohearn1):

Try graphing x^3, x^4, and x^5 and x^6 and see what you notice

OpenStudy (daniel.ohearn1):

The question is to what value n must we take x so we have two solutions Where c = x^n

OpenStudy (daniel.ohearn1):

That's two possible x values that give the same y value, no more, no less

OpenStudy (taylor.joe):

Im back. I apologize.

OpenStudy (daniel.ohearn1):

What did you find about the function of x and you raise it to different values for n?

OpenStudy (taylor.joe):

That there would only be two values and it would be equal to y?

OpenStudy (daniel.ohearn1):

What is y ? x^n=c is the equation you are dealing with

OpenStudy (taylor.joe):

I thought i saw y up there. I must have read something wrong.

OpenStudy (daniel.ohearn1):

y is what you would typically if in your calculator if you wanted to graph x^5 for instance

OpenStudy (taylor.joe):

Ohhhhh okay